Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Kalpan probability question

Problem Solving — algebra and arithmetic
Expert replies
by priyasaibaba » Fri Jun 18, 2010 8:44 pm
Can anyone help me solve this question?

The events A and B are independent. The probability that both events A and B occur is 0.21. The probability that event A occurs and event B does not occur is 0.49. What is the probability that at least one of the events A and B occur?

Can you also explain me the following? This statement was given in Kaplan explanation.

Now when the event A occurs, the event B either occurs or does not occur. So the probability that event A occurs is equal to the probability that both events A and B occur, plus the probability that event A occurs and event B does not occur.

I thought the probability for event A to occur is, Probability that event A only occurs + Probability that both events occurs.

Can someone clarify the above explanation given by Kaplan.

Thanks,
Priya
Join the discussion
Source: — Quantitative Reasoning |

by amising6 » Fri Jun 18, 2010 9:39 pm
priyasaibaba wrote:Can anyone help me solve this question?

The events A and B are independent.
( And B are not depending on each other)The probability that both events A and B occur is 0.21)p(a+B). The probability that event A occurs and event B does not occur is 0.49. What is the probability that at least one of the events A and B occur?


Can you also explain me the following? This statement was given in Kaplan explanation.

Now when the event A occurs, the event B either occurs or does not occur (reason being A is independent of B ,so if A occurs B can occur or cannot occcur). So the probability that event A occurs is equal to the probability that both events A and B occur, plus the probability that event A occurs and event B does not occur.( here they are taking about at least one of the events A and B occur)


I thought the probability for event A to occur is, Probability that event A only occurs ( you can not say only A as it is independent of B)+ Probability that both events occurs.


Can someone clarify the above explanation given by Kaplan.

Thanks,
Priya
yes
Ideation without execution is delusion
Join the discussion

by KapTeacherEli » Mon Jun 21, 2010 11:44 am
Hi Priya,

The general principle of probabilities is that mutually exclusive probabilities add together. In other words, if I am picking a card out of a deck, the probability of getting a King is 4/52 = 1/13 (4 kings our of 52 cards). The probability of picking a Queen is the same. Thus, the probability of a single pick being a King or a Queen is 1/13 + 1/13 = 2/13; I cannot pick a King and a Queen at the same time, so I add the mutually exclusive probabilities. This matches what I would get using the standard probability formula; there are a total of 8 Kings and Queens in the deck, and 8/52 = 2/13.

Here, if event A happens, there are two mutually exclusive possibilities: either B happens or it doesn't. Thus, the odds of A happening and B happening and the odds of A happening and B not happening can be added; the result, 0.6, is the odds of A occurring and B either occurring or not.

Finally, we apply a little common sense; B occurring or not occurring are the only two possibilities, since the default assumption is that events can half-occur. Thus, since we've added the chances of A happening with B and A happening without B, we have accounted for every possible occurrence of A. That leaves us with 0.6 as the final value for the odds of A happening.
Eli Meyer
Kaplan GMAT Teacher
Cambridge, MA
www.kaptest.com/gmat

ImageImageImage
Join the discussion

by ajayiitr » Mon Jun 21, 2010 7:48 pm
The events A and B are independent. The probability that both events A and B occur is 0.21. The probability that event A occurs and event B does not occur is 0.49. What is the probability that at least one of the events A and B occur?

A and B are independent = probability of A happening is not affected by B happening, similarly probability of B happening is not affected by A happening. If on other hand, if probability of A or B is dependent on other, that will be conditional probability. May be GMAT does not test that.

However, they can occur together or not. That will be determined by whether they are mutually exclusive or not. If mutually exclusive, P(A and B) = 0, which is not situation here, as,

P(A and B) = 0.21 = P(A) X P(B)
The probability that event A occurs and event B does not occur is 0.49 implies: P(A=true and B = false) = P(A) X(1-P(B)) = 0.49

Question is: What is the probability that at least one of the events A and B occur? implies: P(A or B) = P(A) + P(B)- P(A and B) = P(A) + P(B) - P(A) X P(B)

Solving from given two eqns, my answer is 0.79
Join the discussion

by gmat1011 » Sun Jun 27, 2010 1:07 am
This formula triggers some memories (I remember this from math class in school!):

P(A or B) = P(A) + P(B)- P(A and B) = P(A) + P(B) - P(A) X P(B)

Can someone clarify if P(A or B) is the same as saying: "probability that at least one of the events A and B occur"

The problem for applying this formula is figuring out the p(B) --- which "ajayiitr" derived above based on p(A) which is given

You get the same result of .79 if you draw a venn diagram - put .21 in the overlapping region and put .49 (for the remaining part of circle A); the remaining part of circle B other than overlapping region is .3 (as .3 + .49 + .21 make 1).

Add .49 + .3 + = .79

Can anyone provide the correct answer - the Venn diagram method seems a bit simplistic and I am not sure if its correctly applied here...
Join the discussion

by dku123 » Wed Jun 30, 2010 4:32 am
amising6 wrote:
priyasaibaba wrote:Can anyone help me solve this question?

The events A and B are independent.
( And B are not depending on each other)The probability that both events A and B occur is 0.21)p(a+B). The probability that event A occurs and event B does not occur is 0.49. What is the probability that at least one of the events A and B occur?


Can you also explain me the following? This statement was given in Kaplan explanation.

Now when the event A occurs, the event B either occurs or does not occur (reason being A is independent of B ,so if A occurs B can occur or cannot occcur). So the probability that event A occurs is equal to the probability that both events A and B occur, plus the probability that event A occurs and event B does not occur.( here they are taking about at least one of the events A and B occur)


I thought the probability for event A to occur is, Probability that event A only occurs ( you can not say only A as it is independent of B)+ Probability that both events occurs.


Can someone clarify the above explanation given by Kaplan.

Thanks,
Priya
yes
This problem can also be solved by using 2*2 matrix(table)
A notA Total

B .21 .09 .30

notB .49 .21 .70

Total .70 .30 1

The probability that at least one event occurs= 1-the probability that neither of them occurs=1-.21=.79 OR
P(A OR B)=P(A)+P(B)-P(AB)=.70+.30-.21=.79,
I think this is helpful and much easier.
Kind Regards
Dinesh
Join the discussion

by mparakala » Tue Feb 05, 2013 11:53 am
Lets denote A and B as follows:

probability of "A" OCCURRING = a
probability of "A" NOT OCCURRING = (1-a)

similarly, B occurring = b
B NOT occurring = (1-b)

given, ab = 0.21;

a * (1-b) = 0.49
=> a-ab = 0.49 (substitute ab=0.21)

a = 0.70 ; (1-a) = 0.30
b = 0.30 ; (1-b) = 0.70

ATLEAST one of A and B occurring = 1 - (Neither A nor B occurring)
= 1 - ([1-a] * [1-b])
= 1- (0.3*0.7)
= 1-0.21
Answer = 0.79
Join the discussion

Re: Kalpan probability question

by Scott@TargetTestPrep » Thu Jul 02, 2020 3:26 pm
priyasaibaba wrote: ↑
Fri Jun 18, 2010 8:44 pm
Can anyone help me solve this question?

The events A and B are independent. The probability that both events A and B occur is 0.21. The probability that event A occurs and event B does not occur is 0.49. What is the probability that at least one of the events A and B occur?

Can you also explain me the following? This statement was given in Kaplan explanation.

Now when the event A occurs, the event B either occurs or does not occur. So the probability that event A occurs is equal to the probability that both events A and B occur, plus the probability that event A occurs and event B does not occur.

I thought the probability for event A to occur is, Probability that event A only occurs + Probability that both events occurs.

Can someone clarify the above explanation given by Kaplan.

Thanks,
Priya
Solution:

The probability that at least one of the events A and B occur is the same as saying the probability that A or B occurs. Similarly, the probability that event A occurs and event B does not occur is the same as saying the probability only A occurs. Recall that:

P(A or B) = P(A) + P(B) - P(A and B)

and

P(A only) = P(A) - P(A and B)

and if A and B are independent, as in this case, we have:

P(A and B) = P(A) x P(B)

From the second equation above, we have:

0.49 = P(A) - 0.21

0.7 = P(A)

From the third equation, we have:

0.21 = 0.7 x P(B)

0.3 = P(B)

Finally, using the first equation, we have:

P(A or B) = 0.7 + 0.3 - 0.21

P(A or B) = 0.79

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion